OpenStudy (turingtest):

(supposedly) Elementary probability: A conservative design team, call it C, and an innovative design team, call it N, are asked to separately design a new product within a month. We know from past experience that: (a) The probability that team C is successful is 2/3 (b) The probability that team N is successful is 1/2 (c) The probability that at least one team is successful is 3/4 Assume that exactly one successful design is produced. What is the probability that it was produced by team N?

OpenStudy (turingtest):

I already have the answer form the book, I just want to see other approaches.

OpenStudy (experimentx):

is answer 1/12 ?

OpenStudy (turingtest):

no, should I say what it is?

OpenStudy (turingtest):

gotta love how this is the first problem in the book... MIT lol

OpenStudy (experimentx):

looks like i have to use conditional probability.

OpenStudy (turingtest):

yeah

OpenStudy (turingtest):

it's in the section on conditional probability, but their solution hardly seems to involve the formula they use

OpenStudy (turingtest):

I mean, their solution doesn't involve the actual formula for conditional probability, which is how I tried (and failed) to solve it. Hence I want to see another approach.

OpenStudy (experimentx):

by that formula ... i seem to get 6/4>1

OpenStudy (turingtest):

I got something>1 when I tried to use the formula as well

OpenStudy (experimentx):

last guess is it 1/24 ?

OpenStudy (turingtest):

no

OpenStudy (experimentx):

looks like i am out of guess. damn probability ... never been my forte.

OpenStudy (turingtest):

should I write their solution, or just the answer, or neither and just try to remember how I tried to work it. This is my first time playing with probability. I like it, but it sure is sneaky.

OpenStudy (uri):

probability is indeed an interesting topic :p

OpenStudy (experimentx):

(a) The probability that team C is successful is 2/3 = P(C) (b) The probability that team N is successful is 1/2 = P(N) (c) The probability that at least one team is successful is 3/4 = P(CUN) using inclusion exclusion i seem to get \( P(C \cap N) = 5/12 \)

OpenStudy (experimentx):

another guess ... is it 1/3 ?

OpenStudy (turingtest):

...sorry, but still no

OpenStudy (turingtest):

It frustrates me that in their solution they don't even speak of \(P(N\cap C),~P(N\cup C)\) or even P(N) or P(C)

OpenStudy (experimentx):

woops!! 1/3 wouldn't work

OpenStudy (turingtest):

If I were to try to model the problem using the formula I am setting up\[P(N|\{N^cC, C^cN\})\]

OpenStudy (turingtest):

I got 2/9 this time.. which is still wrong.

OpenStudy (experimentx):

what is the final answer?

OpenStudy (turingtest):

actually using my method I got 5/6 the actual answer is 1/4

OpenStudy (turingtest):

should I show their working?

OpenStudy (experimentx):

no

OpenStudy (anonymous):

The probability of at least one team being successful is 1 minus the probability that both teams fail, 1 - (1/3)(1/2)= 5/6.

OpenStudy (turingtest):

you didn't make use of the probability of either being successful as 3/4

OpenStudy (anonymous):

Use the odds ratio to get the probability that it was one team rather than the other.

OpenStudy (anonymous):

I did not use that probability because I think it was mistaken. If the probabilities are independent, then 1 - both fail = at least one succeeds.

OpenStudy (turingtest):

how can it be mistaken? it's a given in the problem.

OpenStudy (turingtest):

even if I ignore that, I don't see how your method will give the right answer

OpenStudy (anonymous):

How indeed? ! Sometimes the problems are wrong. How can it be correct? Are they partly collaborating How not independent? Perhaps we are not getting the correct answer because the statement was wrong but their solution was right.

OpenStudy (turingtest):

well, keep in mind this is the first worked problem in a book that has been used for 40-ish years at MIT, so I highly doubt they are mistaken about anything.

OpenStudy (anonymous):

If we have the probability that one team got it right, then the odds favor the team that is more often successful by the ration of the two success rates. "At least one" is the joker here, and my odds ratio may fail for this reason.

OpenStudy (anonymous):

I grant you that it is unlikely that the error has remained for over 40 years.

OpenStudy (turingtest):

1 - both fail = at least one succeeds=5/6=P(N C)+P(N ~C) +P(~N C) how would you propose to wrestle P(N ~C), which is what we want, out of this?

OpenStudy (anonymous):

Given that one was successful, the odds ratio of its being done by N seems to be (1/2)/(2/3) = 3/4 and: 1 and the probability is (3/4)/((1+3/4) = 0.43. x:y odds ratio is x/(x+y) probability.

OpenStudy (turingtest):

I guess I just don't follow, but it still doesn't yield 1/4

OpenStudy (anonymous):

I am probably wrong.

random231 (random231):

its bayes theorem

random231 (random231):

lemme post the question again so i dont hafta scroll up again

random231 (random231):

(supposedly) Elementary probability: A conservative design team, call it C, and an innovative design team, call it N, are asked to separately design a new product within a month. We know from past experience that: (a) The probability that team C is successful is 2/3 (b) The probability that team N is successful is 1/2 (c) The probability that at least one team is successful is 3/4 Assume that exactly one successful design is produced. What is the probability that it was produced by team N?

OpenStudy (anonymous):

Bayes theorem sounds plausible, P(this|that).

OpenStudy (turingtest):

we are nowhere near Baye's theorem in the book yet, but if it works I'd like to to see it.

random231 (random231):

lemme try

random231 (random231):

is the answer 3/7?

OpenStudy (turingtest):

no, I wrote the answer above a few times if you want it

OpenStudy (anonymous):

look at the link I posted

OpenStudy (turingtest):

what about it?

OpenStudy (anonymous):

It is just showing some ideas of how to arrive at the answer

OpenStudy (turingtest):

explain

OpenStudy (turingtest):

seems quite irrelevant to me, what part specifically helps in this problem?

OpenStudy (turingtest):

that's my book ....

OpenStudy (turingtest):

so that's the solution I *don't* want

OpenStudy (anonymous):

lol..it shows how to get the answer

OpenStudy (turingtest):

yes, I'm quite aware of that

OpenStudy (anonymous):

sorry...I tried

OpenStudy (turingtest):

sorry I wanted to delete that so others don't read it... which I'm sure a few have by now

OpenStudy (anonymous):

my bad....I didn't know

OpenStudy (turingtest):

I know, it's cool I think I'll repost this Q, this thread is too long anyway

OpenStudy (anonymous):

Afterthought: I see I mistakenly equated the idea that the teams worked independently to the probabilities' being independent, no overlap P(N v C)=0, when their talents and training could/did mean significant overlap.